OpenStudy (pottersheep):

I need Trig Identity help please..! (Grade 12) How do I prove... (cos2x/1+sin2x) = tan(pi/4- x)

OpenStudy (tkhunny):

You have written this: \(\dfrac{\cos(2x)}{1} + \sin(2x) = \tan(\frac{\pi}{4} - x)\). Is this what you intend?

OpenStudy (pottersheep):

Ohhhhh you are right my mistake!

OpenStudy (pottersheep):

It should be (cos2x/(1+sin2x) = tan(pi/4- x)

OpenStudy (anonymous):

\[\frac{\cos^2(x)}{1+\sin^2(x)}\]?

OpenStudy (pottersheep):

No, They are not to the power of 2, they are just 2x :)

OpenStudy (tkhunny):

Well, it certainly looks like an exercise in double angles. Perhaps expanding all three expressions will lead to something.

OpenStudy (pottersheep):

hmm

OpenStudy (pottersheep):

im still stuck :(

OpenStudy (pottersheep):

I THINK I GOT IT :)

OpenStudy (pottersheep):

yay double angles worked~~~!!

OpenStudy (tkhunny):

Let's see your work and perhaps we can untangle it. \(\cos(2x) = \cos^{2}(x) - \sin^{2}(x) = 1 - 2\sin^{2}(x) = 2\cos^{2}(x) - 1\) \(\sin(2x) = 2\sin(x)\cos(x)\) You may wish to convert the tangent to sine/cosine. Really, the idea behind these things is to get you to EXPLORE these relationships. Don't expect to see a clear solution right away. Play with it until somthing pops out.

OpenStudy (sirm3d):

@pottersheep do you need hint?